OpenStudy (anonymous):

log6 (x+4)=2 solve it with no x terms. give the excat answer

OpenStudy (anonymous):

x+4 = 6^2 x= 32

OpenStudy (anonymous):

logb(x)=y goes to the form b^y = x so he is right, this is just a more general form for if you have more similar problems